# Online Item Cold-Start Recommendation with Popularity-Aware Meta-Learning

### v1

@@ Abstract.
With the rise of e-commerce and short videos, online recommender systems that can capture users’ interests and update new items in real-time play an increasingly important role. In both online and offline recommendation, the cold-start problem due to interaction sparsity has been affecting the recommendation effect of cold-start items, which is also known as the long-tail problem of item distribution. Many cold-start scheme based on fine-tuning or knowledge transferring shows excellent performance on offline recommendation. Yet, these schemes are infeasible for online recommendation on streaming data pipelines due to different training method, computational overhead and time constraints.
Inspired by the above questions, we propose a model-agnostic recommendation algorithm called Popularity-Aware Meta-learning (PAM), to address the item cold-start problem under streaming data settings. PAM divides the incoming data into different meta-learning tasks by predefined item popularity thresholds. The model can distinguish and reweight behavior-related features and content-related features in each task based on their different roles in different popularity levels, thus adapting to recommendations for cold-start samples. These task-fixing design significantly reduces additional computation and storage costs compared to offline methods. Furthermore, PAM also introduced data augmentation and an additional self-supervised loss specifically designed for low-popularity tasks, leveraging insights from high-popularity samples. This approach effectively mitigates the issue of inadequate supervision due to the scarcity of cold-start samples. Experimental results across multiple public datasets demonstrate the superiority of our approach over other baseline methods in addressing cold-start challenges in online streaming data scenarios.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining ; August 3-7, 2025; Toronto, Canadaisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Platforms that depends on online recommendation service, including short video, e-commerce, and streaming media play an indispensable role in people’s lives. Online recommender systems need to capture interest shifts in the system through periodic updates while taking into account the historical information of users and items. Many previous works (Devooght et al., 2015; Zhang et al., 2020; Wang et al., 2020) have demonstrated superior performance in the realm of online recommendation. These online systems are characterized by a large amount of data, strict time-consuming constraints, changes in global information (e.g. user interests, item popularity) over time (Peng et al., 2021) and the requirement of one epoch streaming training. A common way to train the system streamingly is continuously optimizing the model to capture information shifts. The same network parameters are applied for all users and items, and the parameters are updated using data arriving during the period to avoid being stale.
@@
Such cold-start problem has garnered considerable attention and has been addressed through various solutions in recommendation scenarios. The representation-based approach (Lee et al., 2019; Kang et al., 2019; Zhu et al., 2021) considers items as tasks and generates personalized parameters for different items to enhance the effectiveness of cold-start, such as meta-learning (Lee et al., 2019) and knowledge transfer (Zhu et al., 2021). The side-information-based approach (Pan et al., 2022; Yu et al., 2018) focuses on using side information and pay less attention to the item’s popularity-related information, which also have different importance in the recommendation of cold and popular items. However, the methods are still challenging to apply in online scenarios with streaming data.
@@
larity-Awared Meta-learning framework (PAM) for solving the item cold-start problem under streaming data in online systems. Firstly, we propose a fixed task segmentation with predefined popularity thresholds in the gradient-based meta-learning (Finn et al., 2017). We leverage meta-learning’s ability to share meta-parameters between tasks and generate personalized parameters for different tasks. In this way, we optimise the performance of cold-start tasks without losing the interest information of streaming data, and the personalized parameters also avoid making the system over-fitting popular item data, solving the Matthew cold-start problem in online scenarios. In PAM, tasks of varying popularity levels can share meta-knowledge and utilize popularity-related features and content-related features with different weights. Additionally, the fixed task segmentation enables efficient online inference without fine-tuning on each item, thus adapts to streaming training scenarios while addressing the computational overhead and time constraints.
@@ 2.3. Meta-Learning
Meta-learning (Finn et al., 2017), also known as learning to learn, is an emerging few-shot learning method that seeks to quickly adapt to the corresponding task with a few-shot data to achieve better recommendation results. The meta-learning training process divide the training set into support and query set, the former are used to generate unique parameters and the latter are used to calculate loss for training. Since MeLU (Lee et al., 2019), meta-learning has also demonstrated its applicability in the field of recommender systems due to the high overlap between few-shot learning and cold-start recommendation scenarios (Vartak et al., 2017; Lee et al., 2019; Dong et al., 2020; Lu et al., 2020; Bharadhwaj, 2019; Zheng et al., 2020; Neupane et al., 2022; Maheri et al., 2023; Pang et al., 2022; Wu and Zhou, 2023). In recent work, ClusterSeq (Maheri et al., 2023) preserves secondary user preferences through a sequential recommendation system based on meta-learning clustering. Online recommendation is also another scenario where the training process of meta-learning is applicable because of its characteristic of making recommendation model updates in a temporal order (Zhang et al., 2020; Sun et al., 2021; Peng et al., 2021; Xie et al., 2022; Kim et al., 2022; Du et al., 2019). As an example, SML (Zhang et al., 2020) trains a meta-model to output new parameters based on the parameters of the past moments and the data of the current moment. FORM (Sun et al., 2021) adapts user-specific tasks and leverages the online meta leader to optimize online training performance. While FORM alleviates the time cost of fine tuning for individual users, it continues to face challenges in meeting the real-time demands in streaming data scenarios, such as product or video recommendations.
@@ 4.2. Popularity-Aware Meta-Learning
As the cold-start item samples constitute only a small fraction of the online traffic, their influence on model updating is limited. Consequently, the recommender system may exhibit satisfactory performance for popular items but may perform poorly for cold-start items. However, directly excluding samples from popular items could result in reduced performance for both cold-start and popular items. This is because samples from popular items also contain user interest-related information, and the exclusion leads to an information loss.
@@ 5.1.1. Datasets
We selected three public datasets that containing timestamp to ensure that the characteristic of online streaming data is simulated, including MovieLens (Harper and Konstan, 2015), Yelp and Book (Hou et al., 2024).
@@ 5.6. Online A/B Tests (RQ5)
To evaluate the performance of PAM in an online system, we deployed the PAM in a commercial, billion-user-scale, online recommender system and compared it to the PF baseline. In the online system, we counted the following metrics, including the rate of items appearing in users’ recommendations (Show%), and the ratio of users liking (LTR), commenting (CMTR), and collecting videos (CLTR). The results of the experiments in Table 4 show that the high accuracy of PAM for item recommendation enables items to be better recommended to users who favor the item.

### v2

@@ Abstract.
With the rise of e-commerce and short videos, online recommender systems that can capture users’ interests and update new items in real-time play an increasingly important role. In both online and offline recommendation, the cold-start problem due to interaction sparsity has been affecting the recommendation effect of cold-start items, which is also known as the long-tail problem of item distribution. Many cold-start scheme based on fine-tuning or knowledge transferring shows excellent performance on offline recommendation. Yet, these schemes are infeasible for online recommendation on streaming data pipelines due to different training method, computational overhead and time constraints. Inspired by the above questions, we propose a model-agnostic recommendation algorithm called Popularity-Aware Meta-learning (PAM), to address the item cold-start problem under streaming data settings. PAM divides the incoming data into different meta-learning tasks by predefined item popularity thresholds. The model can distinguish and reweight behavior-related and content-related features in each task based on their different roles in different popularity levels, thus adapting to recommendations for cold-start samples. These task-fixing design significantly reduces additional computation and storage costs compared to offline methods. Furthermore, PAM also introduced data augmentation and an additional self-supervised loss specifically designed for low-popularity tasks, leveraging insights from high-popularity samples. This approach effectively mitigates the issue of inadequate supervision due to the scarcity of cold-start samples. Experimental results across multiple public datasets demonstrate the superiority of our approach over other baseline methods in addressing cold-start challenges in online streaming data scenarios.
@@
Yunze Luo, Jingchi Wang, and Kaigui Bian are affiliated with School of CS, AI Innovation Center, National Engineering Laboratory for Big Data Analysis and Applications, State Key Laboratory of Multimedia Information Processing, Peking University.
copyright: acmlicensedjournalyear: 2025doi: XXXX.XXXXXconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining ; August 3-7, 2025; Toronto, Canadaisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Platforms that depends on online recommendation service, including short video, e-commerce, and streaming media, play an indispensable role in people’s lives. Online recommender systems need to capture interest shifts in the system through periodic updates while taking into account the historical information of users and items. Many previous works (Devooght et al., 2015; Zhang et al., 2020; Wang et al., 2020) have demonstrated superior performance in the realm of online recommendation. These online systems are characterized by a large amount of data, strict time-consuming constraints, changes in global information (e.g. user interests, item popularity) over time (Peng et al., 2021) and the requirement of one epoch streaming training. A common way to train the system streamingly is continuously optimizing the model to capture information shifts. The same network parameters are applied for all users and items, and the parameters are updated using data arriving during the period to avoid being stale.
@@
Such cold-start problem has garnered considerable attention and has been addressed through various solutions in recommendation scenarios. The representation-based approach (Lee et al., 2019; Kang et al., 2019; Zhu et al., 2021) considers items as tasks and generates personalized parameters for different items to enhance the effectiveness of cold-start item recommendation, such as meta-learning (Lee et al., 2019) and knowledge transfer (Zhu et al., 2021; Chen et al., 2023). The side-information-based approach (Pan et al., 2022; Yu et al., 2018; Chen et al., 2024) focuses on using content-based side information and pay less attention to the item’s popularity-related information, which also have different importance in the recommendation of cold and popular items. However, the methods above are still challenging to apply in online scenarios with streaming data.
@@
larity-Awared Meta-learning framework (PAM) for solving the item cold-start problem under streaming data in online systems. Firstly, we propose a fixed task segmentation with predefined popularity thresholds in the gradient-based meta-learning (Finn et al., 2017). We leverage meta-learning’s ability to share meta-parameters between tasks and generate personalized parameters for different tasks. In this way, we optimise the performance of cold-start tasks without losing the interest information of streaming data, and the personalized parameters also avoid making the system over-fitting popular item data, solving the Matthew cold-start problem in online scenarios. In PAM, tasks of varying popularity levels can share meta-knowledge and utilize popularity-related features and content-related features with different weights. Additionally, the fixed task segmentation enables efficient online inference without fine-tuning on each item by storing different tasks’ parameters in advance, thus adapts to streaming training scenarios while addressing the computational overhead and time constraints.
@@ 2.3. Meta-Learning
Meta-learning (Finn et al., 2017), also known as learning to learn, is an emerging few-shot learning method that seeks to quickly adapt to the corresponding task with a few-shot data to achieve better recommendation results. The meta-learning training process divide the training set into support and query set, the former are used to generate unique parameters and the latter are used to calculate loss for training. Since MeLU (Lee et al., 2019), meta-learning has also demonstrated its applicability in the field of recommender systems due to the high overlap between few-shot learning and cold-start recommendation scenarios (Vartak et al., 2017; Lee et al., 2019; Dong et al., 2020; Lu et al., 2020; Bharadhwaj, 2019; Zheng et al., 2020; Neupane et al., 2022; Maheri et al., 2023; Pang et al., 2022; Wu and Zhou, 2023). In recent work, ClusterSeq (Maheri et al., 2023) preserves secondary user preferences through a sequential recommendation system based on meta-learning clustering. Online recommendation is also another scenario where the training process of meta-learning is applicable because of its characteristic of making recommendation model updates in a temporal order (Zhang et al., 2020; Sun et al., 2021; Peng et al., 2021; Xie et al., 2022; Kim et al., 2022; Du et al., 2019). As an example, SML (Zhang et al., 2020) trains a meta-model to output new parameters based on the parameters of the past moments and the data of the current moment. FORM (Sun et al., 2021) adapts user-specific tasks and leverages the online meta leader to optimize online training performance.
@@ 4.2. Popularity-Aware Meta-Learning
As the cold-start item samples constitute only a small fraction of the online traffic, their influence on model updating is limited. Consequently, the recommender system may exhibit satisfactory performance for popular items but may perform poorly for cold-start items. However, directly excluding samples from popular items could result in reduced performance for all items. This is because samples from popular items also contain user interest-related information, and the exclusion leads to an information loss.
@@ 5.1.1. Datasets
We selected three public datasets that containing timestamp to ensure that the characteristic of online streaming data is simulated, including MovieLens (Harper and Konstan, 2015), Yelp and Book (Hou et al., 2024). Details of datasets are shown in  A.
@@ 5.6. Online A/B Tests (RQ5)
To evaluate the performance of PAM in an online system, we deployed the PAM in a commercial, billion-user-scale, online recommender system and compared it to the PF baseline. For company privacy, we don’t report the implementation details and the real performance of the original retrieval stage. Instead, we report the performance gain ratio improved by our approach PAM. In the online system, we counted the following metrics, including the rate of items appearing in users’ recommendations (Show%), and the ratio of users liking (LTR), commenting (CMTR), and collecting videos (CLTR). The results of the experiments in Table 4 show that the high accuracy of PAM for item recommendation enables items to be better recommended to users who favor the item.
@@ 6. Conclusion
Acknowledgements.
This work is partially sponsored by National Key R&D Program of China under Grant 2022YFB3104200, NSFC 62XXXXXX and 62032003.
Appendix A Datasets Settings
We selected three public datasets that containing timestamp information.